# Locality-aware Gaussian Compression for Fast and High-quality Rendering

### v1

@@ 1 Introduction
To achieve high compression performance while maintaining the rendering efficiency of the 3DGS rendering pipeline, in this paper, we present LocoGS, a novel locality-aware compact 3DGS framework. As will be presented in the analysis in Section 4.1, Gaussians exhibit strong local coherence across nearby Gaussians. Inspired by this, our framework introduces a novel 3D Gaussian representation that leverages the spatial coherence of Gaussians. Our representation exploits a grid-based neural field representation that allows compact modeling of continuous physical quantities in volumetric scenes (Chen et al., 2022; Müller et al., 2022; Tang et al., 2022). Specifically, our representation encodes locally-coherent Gaussian attributes into a multi-resolution hash grid (Müller et al., 2022) that enables a sparse voxel-grid representation to characterize the sparsely distributed structure of point-based representation. Unlike anchor-based approaches, our representation allows the usage of the original 3DGS rendering pipeline without any modification, thus maintaining the rendering efficiency of the original pipeline.
@@ 2 Related Work
Neural fields have emerged as a dominant representation for various continuous signals, including images (Dupont et al., 2021; 2022; Strümpler et al., 2022), signed distances (Yu et al., 2022; Wang et al., 2023), and 3D/4D scenes (Mildenhall et al., 2021; Barron et al., 2021; 2022; Müller et al., 2022). In particular, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) have revolutionized the domain of neural rendering, exhibiting a remarkable ability to represent volumetric scenes. Recently, grid-based explicit representations have been adopted to encode local features and exploit them to model 3D scenes (Chen et al., 2022; Sun et al., 2022; Müller et al., 2022; Reiser et al., 2023). Despite the outstanding performance, they yield limitations regarding the heavy storage burden. To address this issue, several approaches exploit efficient grid structure (Chen et al., 2022; Müller et al., 2022; Tang et al., 2022), pruning (Fridovich-Keil et al., 2022; Rho et al., 2023), and quantization (Takikawa et al., 2022; Shin & Park, 2024) to further consider efficiency in in terms of storage. Although compact neural radiance fields significantly reduce storage costs, the persistent issue of slow rendering speed remains a limiting factor to becoming a practical 3D representation. In this paper, we aim to accomplish real-time rendering via efficient point-based rendering while leveraging the compactness of neural fields to attain high rendering quality.
@@
C3DGS (Niedermayr et al., 2024) sorts Gaussian attributes according to their positions along a space-filling curve, placing spatially close Gaussians at nearby positions in the sorted sequence. This enables subsequent compression steps to exploit local coherence. However, projecting 3D Gaussians to a 1D sequence along a space-filling curve cannot fully leverage the local coherence of Gaussian attributes. Compact-3DGS (Lee et al., 2024) uses a hash grid to encode view-dependent colors and vector quantization for geometric attributes such as rotation and scale. While it achieves a noticeable 81.8× compression ratio for color, it only achieves a 5.5× compression ratio for vector-quantized attributes, as it exploits global statistics instead of local coherence.
@@ 6.1 Comparison
We compare our method with existing compact 3D Gaussian frameworks as well as the original 3DGS (Kerbl et al., 2023) and Scaffold-GS Lu et al. (2024). For comparison, we roughly categorize the existing compression approaches into the quantization-based, image-codec-based, hash-grid-based, and anchor-based approaches. The quantization-based approaches include CompGS (Navaneet et al., 2024), C3DGS (Niedermayr et al., 2024), LightGaussian (Fan et al., 2023), and EAGLES (Girish et al., 2023). The image-codec-based approach includes SSGS (Morgenstern et al., 2023), which maps 3D Gaussian attributes onto a 2D plane and compresses the attributes using an image codec. The hash-grid-based approaches include Compact-3DGS (Lee et al., 2024) as well as ours. Finally, the anchor-based approaches include Scaffold-GS (Lu et al., 2024) and HAC (Chen et al., 2024), while Scaffold-GS adopts an anchor-based representation not for compression but for improving the rendering quality. For a fair comparison of rendering quality, storage size, and rendering speed, we reproduce the results of all the baseline models using the officially released codes and configurations. Refer to appendix for details on the rendering speed measurement. In our evaluation, we evaluate two variants of our approach: ‘Ours’, and ‘Ours-Small’. ‘Ours’ is our baseline model, while ‘Ours-Small’ uses a smaller hash grid and more aggressive pruning compared to ‘Ours’ to obtain more compact representations.
@@
Fig. 4, Table 1 and Table 2 show quantitative comparisons of the rendering quality (PSNR, SSIM, LPIPS), storage size, and rendering speed of different approaches. The comparison results show that our approach outperforms the baseline methods, achieving highly compact storage sizes (54.6$\times$ to 96.6$\times$ compressed storage size than 3DGS Kerbl et al. (2023)), while achieving comparable or higher rendering quality. In terms of rendering speed, our approach offers 2.1$\times$ to 2.4$\times$ rendering speed compared to 3DGS. On the other hand, the quantization-based approaches (CompGS (Navaneet et al., 2024), C3DGS (Niedermayr et al., 2024), LightGaussian (Fan et al., 2023), and EAGLES (Girish et al., 2023)) suffer from relatively large storage sizes and lower rendering qualities as they solely rely on global statistics of Gaussian attributes not considering the local coherence. SSGS (Morgenstern et al., 2023), an image-codec-based approach, also exhibits lower compression performance as projecting 3D Gaussians to a 2D image plane cannot fully exploit the local coherence of 3D Gaussians. Compact-3DGS (Lee et al., 2024), which is a hash-grid-based approach, uses a hash grid only as a mapping to an RGB color from a 3D position and a view direction. Thus, it exhibits not only low compression ratios, but also slow rendering speed as its rendering process requires an additional MLP to obtain color values for a given view direction. Finally, HAC (Chen et al., 2024), an anchor-based approach, achieves impressive performance in terms of rendering quality and storage size, but its rendering speed is significantly slower than ours, as its rendering process also requires additional MLPs.
Fig. 5 demonstrates a qualitative comparison of LocoGS against previous methods (Navaneet et al., 2024; Lee et al., 2024; Niedermayr et al., 2024; Fan et al., 2023; Girish et al., 2023; Morgenstern et al., 2023; Chen et al., 2024). Existing compact approaches struggle with the intricate structures of real-world scenes, leading to severe artifacts (e.g., floaters) due to the lack of information for achieving storage efficiency. In contrast, our method clearly shows superior rendering quality, capturing fine details of textureless regions, although it requires minimal storage capacity. Moreover, the result of ‘Ours-Small’, which requires a significantly smaller storage size, surpasses the rendering quality of all the other approaches. Additional quantitative comparisons and qualitative examples can be found in appendix.
@@ 6.2 Ablation Study
As the table shows, locality-aware representation greatly reduces the storage requirement compared to the ‘3DGS + Pruning + Compression’, while maintaining a similar number of Gaussians. The dense initialization further enhances the rendering quality and speed while reducing the storage size as it helps accurately reconstruct the target scene with fewer Gaussians. The adaptive SH bandwidth scheme greatly enhances the rendering speed at the cost of a slightly increased storage size. Furthermore, G-PCC and the other compression steps for the explicit and implicit attributes introduce a significant storage size reduction while maintaining the rendering quality and speed. Firstly, G-PCC decreases $11.7\%$ of storage space by compressing the positions. Moreover, explicit attribute compression saves $15.3\%$ of storage capacity by compressing the base scales and base colors into smaller bits. Finally, implicit attribute compression achieves a further reduction of $41.6\%$ in storage capacity by compressing the parameters of hash-grid and MLPs. Consequently, the compression pipeline shows a 3.18×3.18$\times$3.18 × compression ratio with a minimal loss in rendering performance. appendix provides additional ablation studies on the impact of the dense initialization and adaptive SH bandwidth scheme.

### v2

@@ 1 Introduction
To achieve high compression performance while maintaining the rendering efficiency of the 3DGS rendering pipeline, in this paper, we present LocoGS, a novel locality-aware compact 3DGS framework. As will be presented in the analysis in Section 4.1, Gaussians exhibit strong local coherence across nearby Gaussians. Inspired by this, our framework introduces a novel 3D Gaussian representation that leverages the spatial coherence of Gaussians. Our representation exploits a grid-based neural field representation that allows compact modeling of continuous physical quantities in volumetric scenes (Chen et al., 2022; Müller et al., 2022; Tang et al., 2022). Specifically, our representation encodes locally coherent Gaussian attributes into a multi-resolution hash grid (Müller et al., 2022) that enables a sparse voxel-grid representation to characterize the sparsely distributed structure of point-based representation. Unlike anchor-based approaches, our representation allows the usage of the original 3DGS rendering pipeline without any modification, thus maintaining the rendering efficiency of the original pipeline.
@@ 2 Related Work
Neural fields have emerged as a dominant representation for various continuous signals, including images (Dupont et al., 2021; Strümpler et al., 2022), signed distances (Yu et al., 2022; Wang et al., 2023), and 3D/4D scenes (Mildenhall et al., 2021; Barron et al., 2021; 2022; Müller et al., 2022). In particular, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2021) have revolutionized the domain of neural rendering, exhibiting a remarkable ability to represent volumetric scenes. Recently, grid-based explicit representations have been adopted to encode local features and exploit them to model 3D scenes (Chen et al., 2022; Sun et al., 2022; Müller et al., 2022). Despite the outstanding performance, they yield limitations regarding the heavy storage burden. To address this issue, several approaches exploit efficient grid structure (Chen et al., 2022; Müller et al., 2022), pruning (Fridovich-Keil et al., 2022; Rho et al., 2023), and quantization (Takikawa et al., 2022; Shin & Park, 2024) to further consider efficiency in in terms of storage. Although compact neural radiance fields significantly reduce storage costs, the persistent issue of slow rendering speed remains a limiting factor to becoming a practical 3D representation. In this paper, we aim to accomplish real-time rendering via efficient point-based rendering while leveraging the compactness of neural fields to attain high rendering quality.
@@
C3DGS (Niedermayr et al., 2024) sorts Gaussian attributes according to their positions along a space-filling curve, placing spatially close Gaussians at nearby positions in the sorted sequence. This enables subsequent compression steps to exploit local coherence. However, projecting 3D Gaussians to a 1D sequence along a space-filling curve cannot fully leverage the local coherence of Gaussian attributes. Compact-3DGS (Lee et al., 2024) uses a hash grid to encode view-dependent colors and vector quantization for geometric attributes such as rotation and scale. While it achieves a noticeable 81.8$\times$ compression ratio for color, it only achieves a 5.5$\times$ compression ratio for vector-quantized attributes, as it exploits global statistics instead of local coherence.
@@ 6.1 Comparison
We compare our method with existing compact 3D Gaussian frameworks as well as the original 3DGS (Kerbl et al., 2023) and Scaffold-GS (Lu et al., 2024). For comparison, we roughly categorize the existing compression approaches into the quantization-based, image-codec-based, hash-grid-based, and anchor-based approaches. The quantization-based approaches include CompGS (Navaneet et al., 2024), C3DGS (Niedermayr et al., 2024), LightGaussian (Fan et al., 2023), and EAGLES (Girish et al., 2023). The image-codec-based approach includes SSGS (Morgenstern et al., 2023), which maps 3D Gaussian attributes onto a 2D plane and compresses the attributes using an image codec. The hash-grid-based approaches include Compact-3DGS (Lee et al., 2024) as well as ours. Finally, the anchor-based approaches include Scaffold-GS (Lu et al., 2024) and HAC (Chen et al., 2024), while Scaffold-GS adopts an anchor-based representation not for compression but for improving the rendering quality. For a fair comparison of rendering quality, storage size, and rendering speed, we reproduce the results of all the baseline models using the officially released codes and configurations. In our evaluation, we evaluate two variants of our approach: ‘Ours’, and ‘Ours-Small’. ‘Ours’ is our baseline model, while ‘Ours-Small’ uses a smaller hash grid and more aggressive pruning compared to ‘Ours’ to obtain more compact representations.
@@
Fig. 4, Table 1 and Table 2 show quantitative comparisons of the rendering quality (PSNR, SSIM, LPIPS), storage size, and rendering speed of different approaches. The comparison results show that our approach outperforms the baseline methods, achieving highly compact storage sizes (54.6$\times$ to 96.6$\times$ compressed storage size than 3DGS (Kerbl et al., 2023)), while achieving comparable or higher rendering quality. In terms of rendering speed, our approach offers 2.1$\times$ to 2.4$\times$ rendering speed compared to 3DGS. On the other hand, the quantization-based approaches (CompGS (Navaneet et al., 2024), C3DGS (Niedermayr et al., 2024), LightGaussian (Fan et al., 2023), and EAGLES (Girish et al., 2023)) suffer from relatively large storage sizes and lower rendering qualities as they solely rely on global statistics of Gaussian attributes not considering the local coherence. SSGS (Morgenstern et al., 2023), an image-codec-based approach, also exhibits lower compression performance as projecting 3D Gaussians to a 2D image plane cannot fully exploit the local coherence of 3D Gaussians. Compact-3DGS (Lee et al., 2024), which is a hash-grid-based approach, uses a hash grid only as a mapping to an RGB color from a 3D position and a view direction. Thus, it exhibits not only low compression ratios, but also slow rendering speed as its rendering process requires an additional MLP to obtain color values for a given view direction. Finally, HAC (Chen et al., 2024), an anchor-based approach, achieves impressive performance in terms of rendering quality and storage size, but its rendering speed is significantly slower than ours, as its rendering process also requires additional MLPs.
Fig. 5 demonstrates a qualitative comparison of LocoGS against previous methods (Navaneet et al., 2024; Lee et al., 2024; Niedermayr et al., 2024; Fan et al., 2023; Girish et al., 2023; Morgenstern et al., 2023; Chen et al., 2024). Existing compact approaches struggle with the intricate structures of real-world scenes, leading to severe artifacts (e.g., floaters) due to the lack of information for achieving storage efficiency. In contrast, our method clearly shows superior rendering quality, capturing fine details of textureless regions, although it requires minimal storage capacity. Moreover, the result of ‘Ours-Small’, which requires a significantly smaller storage size, surpasses the rendering quality of all the other approaches.
@@ 6.2 Ablation Study
As the table shows, locality-aware representation greatly reduces the storage requirement compared to the ‘3DGS + Pruning + Compression’, while maintaining a similar number of Gaussians. The dense initialization further enhances the rendering quality and speed while reducing the storage size as it helps accurately reconstruct the target scene with fewer Gaussians. The adaptive SH bandwidth scheme greatly enhances the rendering speed at the cost of a slightly increased storage size. Furthermore, G-PCC and the other compression steps for the explicit and implicit attributes introduce a significant storage size reduction while maintaining the rendering quality and speed. Firstly, G-PCC decreases $11.7\%$ of storage space by compressing the positions. Moreover, explicit attribute compression saves $15.3\%$ of storage capacity by compressing the base scales and base colors into smaller bits. Finally, implicit attribute compression achieves a further reduction of $41.6\%$ in storage capacity by compressing the parameters of hash-grid and MLPs. Consequently, the compression pipeline shows a 3.18×3.18$\times$3.18 × compression ratio with a minimal loss in rendering performance. Appendix provides additional ablation studies on the dense initialization and adaptive SH bandwidth scheme.
@@ 7 Conclusion
8 Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by Korea government (Ministry of Education) (No.2022R1A6A1A03052954, Basic Science Research Program (10%)), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by Korea government (MSIT) (No.RS-2019-II191906, AI Graduate School Program (POSTECH) (10%); No.RS-2021-II211343, AI Graduate School Program (Seoul National University) (5%); No.RS-2024-00457882, AI Research Hub Project (30%); No.RS-2023-00227993, Detailed 3D reconstruction for urban areas from unstructured images (45%)).